An ion engine has a 1T magnetic field and hydrogen plasma is to be short out at an E×B velocity of 1150km/s. how much internal electric field must be present in the plasma?
1
Expert's answer
2020-05-14T09:21:07-0400

As per the question,

Magnetic field of the ion engine "(B)=1T"

Speed of the light "(c)=1150 km/sec=1150\times 10^{3} m/sec"

internal electric field (E)=?

E"=cB = 1150\times 10^{3}\times 1 N/c"

"=1.15\times 10^{6} N/c"